OpenStudy (anonymous):

Split the Function into piece wise defined one f(x) = |x+1| + |x-3| + | | x+5| |

OpenStudy (anonymous):

@shubhamsrg @experimentX

OpenStudy (anonymous):

@mukushla @joemath314159

OpenStudy (anonymous):

Consider\[x<-5\]\[-5<x<-1\]\[-1<x<3\]\[x>3\]

OpenStudy (anonymous):

what to do with that double modulus ?

OpenStudy (anonymous):

nothing to do with that for example for \(x<-5\)\[f(x) = |x+1| + |x-3| + | | x+5| | = |x+1| + |x-3| + | x+5| \]\[f(x)=-(x+1)-(x-3)-(x+5)=-3x-3\]

OpenStudy (anonymous):

i've never seen something like that :D